OpenStudy (anonymous):

Choose the equation below that represents the line passing through the point (1, -4) with a slope of one half. y - 4 = one half(x - 1) y + 4 = 2(x - 1) y + 4 = one half(x - 1) y - 4 = 2(x + 1)

OpenStudy (anonymous):

@Godlovesme

Nnesha (nnesha):

point slope is \[y-y _{1} = m(x-x _{1})\]

Nnesha (nnesha):

replace m by 1/2 and replace y1 and x1 by this order pair (1 , -4)

OpenStudy (godlovesme):

y1 and x1 is like the one we did in the previous m is slope

OpenStudy (anonymous):

y - 4 = 1/2 x - 1 ? I plugged them in , Is that correct?

Nnesha (nnesha):

well noep

Nnesha (nnesha):

there is negative sign in the formula and your y is also negative

OpenStudy (godlovesme):

y- -4 =? @FEARLESS_JOCEY

OpenStudy (anonymous):

Oh, so it's supposed to be a positive 4 not a negative right?

Nnesha (nnesha):

 that's how you should replace